# FaceCoT: A Benchmark Dataset for Face Anti-Spoofing with Chain-of-Thought Reasoning

### v1

@@ B.2 Reinforcement Learning in CoT-Enhanced Progressive Learning (CEPL) framework
Specifically, after completing the two-stage training with CEPL, we introduce a third stage of RL. In this stage, we augment the original multi-task loss—consisting of CoT reasoning and classification supervision—with an additional RL objective driven by our dual-reward functions for semantic accuracy and format compliance. The RL procedure follows the same policy-optimization paradigm described previously, with one key difference: no new data is incorporated. Instead, we directly reuse the image–text pairs employed during the second stage. This design tests whether strategic optimization of output structure and semantics alone—absent any additional training examples—can still yield significant performance gains.
Appendix C Experiments
C.1 The effect of Reinforcement Learning on CoT annotation quality
@@
By introducing RL into the training of the caption model, the annotation accuracy is effectively improved. Although this metric demonstrates the effectiveness of RL in enhancing conclusion label accuracy, relying solely on conclusion accuracy is not sufficient to fully evaluate the semantic quality of the generated annotations. To further verify the advantages of RL-generated CoT annotations in terms of linguistic coherence and logical consistency, we design an experiment in which we treat two sets of generated CoT annotations—those produced by a model trained solely via Supervised Fine-Tuning (SFT) and those produced by an SFT-trained model further refined with RL—as separate training sets under our proposed CoT-Enhanced Progressive Learning (CEPL) framework. The comparative results, reported in Table 8, demonstrate the tangible benefits of RL on CoT data quality: RL not only enhances annotation accuracy, but also significantly improves the consistency, coherence, and semantic reliability of the generated CoT explanations.
C.2 The effect of FaceCoT under low-resolution
@@
In the ablation study presented in the main text, we evaluate the impact of FaceCoT data using an input resolution of $448\times 448448 × 448$. Given that most existing FAS methods [45, 32, 14] conduct experiments at a resolution of $224\times 224224 × 224$, we perform an ablation study at this resolution to verify the effectiveness of our CoT data under low-resolution settings. We compare a single-stage training regime using only label classification data against a single-stage joint training regime incorporating CoT data. As shown in Table 8, the model trained with CoT data achieves a 6.70% reduction in HTER and a 4.61% increase in AUC. These findings can be summarized as follows: (1) The relative gain at $224\times 224224 × 224$ (–5.79% HTER, +3.63% AUC) is substantially larger than at $448\times 448448 × 448$ (–1.42% HTER, +1.54% AUC), demonstrating that our CoT annotations help the model recover fine-grained facial cues that are otherwise lost at lower resolutions. (2) Even when applied in a simple single-stage joint training regime, the CoT-augmented model already outperforms current state-of-the-art methods, demonstrating its superior generalization and robustness conferred by CoT data training.
@@
Table 8: Ablation study on CoT data at a resolution of $224\times 224224 × 224$, with comparison to a setup using only binary label data.
@@
Appendix D Limitation
Although FaceCoT achieves significant advances in annotation scale and logical structure, it has a limitation: it is primarily built upon WFAS and CelebA-Spoof—two of the largest publicly available FAS datasets—it remains inherently challenging to comprehensively capture the full diversity of real-world capture devices and environmental conditions.
Appendix E Demonstration of result interpretability
@@
(a) An interpretable result case on CASIA-MFSD
(b) An interpretable result case on CASIA-SURF-3DMask

### v2

@@ 2.2 Reinforcement Learning in CoT-Enhanced Progressive Learning (CEPL)
Specifically, after completing the two‐stage training with CEPL, we introduce a third stage of RL. In this stage, we augment the original multi-task loss, which consists of CoT reasoning and classification supervision, with an additional RL objective driven by our dual-reward functions for semantic accuracy and format compliance. The RL procedure follows the same policy‐optimization paradigm described previously, with one key difference: no new data is incorporated. Instead, we directly reuse the image–text pairs employed during the second stage. This design tests whether strategic optimization of output structure and semantics, without any additional training examples, can still yield significant performance gains.
2.3 Details for Evaluation Metrics
Since standard FAS metrics such as AUC and HTER require continuous confidence scores rather than binary predictions, we adapt the output of VLMs to provide probabilistic scores. Specifically:
  1. 1.
    Deterministic decoding. To ensure output consistency and avoid randomness from beam search, we set the generation beam number to 1.
  2. 2.
    Token-level logits extraction. Instead of directly treating textual outputs (e.g., yes vs. no) as hard labels, we extract token-level logits from the first generated token. In particular, we identify the token IDs corresponding to ’Yes’ and ’No.’
  3. 3.
    Probability computation. We compute the softmax probability over the two logits, obtaining the confidence that a sample is real:
    $p_{\text{real}}=\frac{\exp(\ell_{\text{No}})}{\exp(\ell_{\text{No}})+\exp(\ell_{\text{Yes}})},$ (2)
    where $\ell_{\text{Yes}}$ and $\ell_{\text{No}}$ denote the logits of the “Yes” and “No” tokens, respectively.
The resulting probability $p_{\text{real}}$ is then used to calculate AUC and HTER following standard definitions in the FAS literature. This procedure allows us to fairly evaluate LLM-based classifiers under conventional spoofing metrics.
3 Experiments
3.1 Cross-Domain Generalization under Widely Adopted Protocol
In the FAS literature, a widely adopted evaluation protocol is the leave-one-out cross-domain testing on four benchmarks: OULU-NPU (O) [3], CASIA-MFSD (C) [46], Idiap Replay-Attack (I) [6], and MSU-MFSD (M) [38]. However, the performance under this protocol has already saturated (with AUC exceeding 99%), making it less discriminative for assessing fine-grained improvements. Therefore, in the main text we focus on a more challenging and generalization-oriented one-to-eleven protocol, which better highlights the advantages of our method. Nevertheless, to further demonstrate the robustness of our approach, we also report results under the widely used leave-one-out protocol. Specifically, we first apply our FAS caption model to generate CoT annotations for the training splits of the O, C, M, and I datasets. Based on these annotated datasets, we then perform four cross-dataset evaluations following the standard protocols. For example, the protocol OCI→M denotes that the model is trained on OULU-NPU, CASIA-MFSD, and Idiap Replay-Attack, and tested on MSU-MFSD. Similarly, OMI→C, OCM→I, and ICM→O are defined in the same manner. As shown in Table 7, our method outperforms previous state-of-the-art methods, achieving the best average HTER and AUC. These results confirm the effectiveness of our approach in improving generalization in FAS.
Table 7: Cross-dataset evaluation results under widely used cross-domain protocol.
[…]
HTER(%) AUC(%) HTER(%) AUC(%) HTER(%) AUC(%) HTER(%) AUC(%) HTER(%) AUC(%)
FGHV [21] 9.17 96.92 12.47 93.47 16.29 90.11 13.58 93.55 12.88 93.51
GDA [51] 9.20 98.00 12.20 93.00 10.00 96.00 14.40 92.60 11.45 94.90
PatchNet [33] 7.10 98.46 11.33 94.58 13.40 95.67 11.82 95.07 10.91 95.95
SSAN [36] 6.67 98.75 10.00 96.67 8.88 96.79 13.72 93.63 9.82 96.46
IADG [49] 5.41 98.19 8.70 96.40 10.62 94.50 8.86 97.14 8.40 96.56
UDG-FAS [25] 5.95 98.47 9.82 96.76 5.86 98.62 10.97 95.36 8.15 97.30
TTDG [48] 4.16 98.48 7.59 98.18 9.62 98.18 10.00 96.15 7.84 97.75
SA-FAS [32] 5.95 96.55 8.78 95.37 6.58 97.54 10.00 96.23 7.83 96.42
DIVT-M [18] 2.86 99.14 8.67 96.92 3.71 99.29 13.06 94.04 7.08 97.35
GAC-FAS [15] 5.00 97.56 8.20 95.16 4.29 98.87 8.60 97.16 6.52 97.19
FLIP [31] 4.95 98.11 0.54 99.98 4.25 99.07 2.31 99.63 3.01 99.20
CFPL [19] 1.43 99.28 2.56 99.10 5.43 98.41 2.50 99.42 2.98 99.05
I-FAS [43] 0.32 99.88 0.04 99.99 3.22 98.48 1.74 99.66 1.33 99.50
Ours 0.42 99.92 0.00 100.00 1.00 99.83 2.81 99.63 1.06 99.85
Table 8: Performance comparison across different spoof types in the Rose-Youtu test set before and after SFT with FaceCoT.
Cheat Type Meaning Number Acc. (Zero-shot) Acc. (SFT) Change
Mc Mask: Cut eyes & mouth 202 100.00% 100.00%
Mf Mask: Full face 100 74.00% 100.00% +26.00%
Mu Mask: Upper part cut 198 93.43% 100.00% +6.57%
Pq Printed paper (quivering) 200 0.00% 95.50% +95.50%
Ps Printed paper (still) 200 0.00% 68.00% +68.00%
Vl Video (Lenovo LCD) 201 0.00% 96.02% +96.02%
Vm Video (Mac LCD) 199 0.00% 71.36% +71.36%
3.2 Fine-grained Analysis on Spoof Type Robustness
To examine whether FaceCoT introduces bias toward certain spoof types, we conduct a fine-grained analysis on the Rose-Youtu [16], which contains seven representative spoofing attack types. We report the per-type detection accuracy before and after fine-tuning with FaceCoT. As shown in Table 8, our approach achieves consistent improvements across all attack categories. These results indicate that FaceCoT provides more comprehensive semantic supervision and enhances general spoof detection capability, rather than overfitting to the dominant categories in the training set.
3.3 Comparison of Zero-shot and CoT-trained Models
To assess the effectiveness of our supervised CoT training, we first establish a zero-shot baseline, where large vision-language models are directly prompted with natural language to classify real versus spoof images without any fine-tuning. To further validate the robustness of our approach, we also conduct the same experiment with Qwen2.5-VL [1], a recent advanced multimodal VLM. As shown in Table 9, both VLMs perform worse than the SOTA method I-FAS [43] under the zero-shot setting. After applying our CoT-based fine-tuning, we observe consistent and substantial gains on both models: MiniCPMV achieves a reduction of 11.61% in HTER and an improvement of 10.45% in AUC, while Qwen2.5-VL shows similar improvements (HTER reduced by 10.28% and AUC improved by 11.97%). These results demonstrate that our FaceCoT dataset and CEPL framework provide stable and significant benefits across different VLM architectures, enabling stronger discriminative ability and cross-domain generalization than relying on zero-shot reasoning alone.
Table 9: Comparison between zero-shot baselines and our supervised CoT method across different backbone VLM models.
Method Avg. HTER (%) Avg. AUC (%)
I-FAS [43] 11.31 93.71
Zero-shot(Qwen2.5-VL-7B [1]) 19.60 83.75
Zero-shot(MiniCPMV-2.6-8B [40]) 17.91 87.32
Ours(Qwen2.5-VL-7B [1]) 9.32 ($\downarrow$10.28) 95.72 ($\uparrow$11.97)
Ours(MiniCPMV-2.6-8B) 6.30 ($\downarrow$11.61) 97.77 ($\uparrow$10.45)
3.4 The effect of Reinforcement Learning on CoT annotation quality
@@
By introducing RL into the training of the caption model, the annotation accuracy is effectively improved. Although this metric demonstrates the effectiveness of RL in enhancing conclusion label accuracy, relying solely on conclusion accuracy is not sufficient to fully evaluate the semantic quality of the generated annotations. To further verify the advantages of RL-generated CoT annotations in terms of linguistic coherence and logical consistency, we design an experiment in which we treat two sets of generated CoT annotations—those produced by a model trained solely via Supervised Fine‐Tuning (SFT) and those produced by an SFT‐trained model further refined with RL—as separate training sets under our proposed CoT‐Enhanced Progressive Learning (CEPL) framework. The comparative results, reported in Table 11, demonstrate the tangible benefits of RL on CoT data quality: RL not only enhances annotation accuracy, but also significantly improves the consistency, coherence, and semantic reliability of the generated CoT explanations.
3.5 The effect of FaceCoT under low-resolution
@@
In the ablation study presented in the main text, we evaluate the impact of FaceCoT data using an input resolution of $448\times 448$. Given that most existing FAS methods [43, 31, 14] conduct experiments at a resolution of $224\times 224$, we perform an ablation study at this resolution to verify the effectiveness of our CoT data under low-resolution settings. We compare a single‐stage training regime using only label classification data against a single‐stage joint training regime incorporating CoT data. As shown in Table 11, the model trained with CoT data achieves a 6.70% reduction in HTER and a 4.61% increase in AUC. These findings can be summarized as follows: (1) The relative gain at $224\times 224$ (–5.79% HTER, +3.63% AUC) is substantially larger than at $448\times 448$ (–1.42% HTER, +1.54% AUC), demonstrating that our CoT annotations help the model recover fine‐grained facial cues that are otherwise lost at lower resolutions. (2) Even when applied in a simple single‐stage joint training regime, the CoT‐augmented model already outperforms current state‐of‐the‐art methods, demonstrating its superior generalization and robustness conferred by CoT data training.
@@
Table 11: Ablation study on CoT data at a resolution of $224\times 224$, with comparison to a setup using only binary label data.
@@
4 Usage of LLMs
In this work, Large Language Models (LLMs) were employed as auxiliary tools in two aspects:
  • Manuscript Refinement: LLMs were used to assist in language polishing and grammar checking after the human authors had completed the technical writing. The scientific content, experiment design, and analysis were fully conducted by the authors.
  • Annotation of FaceCoT-Gold100K: GPT-4o was used to generate Chain-of-Thought (CoT) annotations. Specifically, we carefully designed prompts to guide the model toward describing spoof-related visual cues (e.g., reflection artifacts, cutting marks), rather than allowing free-form generation. The model outputs were subsequently reviewed and refined by human experts to ensure accuracy, fairness, and domain relevance.
5 Limitation and feature work
We have included as many spoofing types as practically possible in FaceCoT to ensure broad coverage. Some uncommon real-world variations in devices or environments are inevitably not captured, which we leave for future extension. Furthermore, while our work primarily emphasizes the utility of FaceCoT for model training and performance gains on downstream FAS tasks, we acknowledge that its potential as a standardized benchmark for evaluation has not been fully explored. In particular, since FaceCoT provides reasoning chains, it can serve as a valuable resource to assess not only predictive accuracy but also the interpretability and coherence of model outputs. We leave a more systematic investigation of FaceCoT’s role in model evaluation and benchmarking as an important direction for future work.
6 Ethics statement
In designing FaceCoT, attention was paid to fairness, bias mitigation, and data privacy, ensuring that the dataset not only enhances interpretability and generalization in Face Anti-Spoofing (FAS), but also adheres to responsible research principles. To address these potential concerns, this section outlines our efforts in three dimensions: dataset fairness, language model bias mitigation, and data privacy protection.
6.1 Dataset Bias and Fairness
The original FAS datasets(e.g., CelebA-Spoof [45] and WFAS [34]) used in our work were not explicitly designed with fairness auditing or demographic balance in mind. The goal of FaceCoT, however, is to introduce a reasoning-based multimodal framework that improves interpretability and generalization in FAS models. This also provides a structured avenue for detecting and mitigating bias through Chain-of-Thought (CoT) rationales. To align FaceCoT with responsible research practices and address potential fairness concerns, we incorporated the following safeguards during our data collection and annotation process:
  • No New Image Data Introduced: All FaceCoT annotations are derived exclusively from publicly available FAS datasets. We did not collect or distribute any new images, and the released dataset contains only annotations indexed to existing data.
  • Bias-Aware Annotation Pipeline: (1) Prompt Design: CoT generation prompts were carefully crafted to steer the model toward spoof-specific visual cues (e.g., reflection artifacts and cutting marks), while explicitly excluding references to race, gender, age, or identity. (2) Human-in-the-Loop Filtering: Expert annotators reviewed all FaceCoT-Gold100K outputs and were instructed to remove any content with identity-based or stereotypical language. (3) Random Auditing: Manual sampling and inspection of 5,000 annotations from FaceCoT-Silver982K revealed no evidence of demographic bias or stereotype leakage. (4) Model Validation Across Subgroups: Models trained with FaceCoT demonstrated balanced performance across evaluation datasets, including with respect to skin tone and gender, indicating no observable subgroup bias attributable to the annotations. As shown in the outputs across various evaluation datasets in our Appendix 7, the model focused on spoofing features related to FAS rather than attributes such as age, gender, or skin tone.
6.2 Use of GPT-4o and Mitigation of Language Model Bias
Large foundation models such as GPT-4o may introduce bias. To mitigate such risks, we employed several safeguards:
  • Constrained Use: GPT-4o was used solely to generate CoT explanations under strict prompt constraints and was never involved in classification or decision-making tasks.
  • Annotation Safeguards: (1) FaceCoT-Gold100K annotations were reviewed and refined by human experts to eliminate any inappropriate or biased language. (2) For FaceCoT-Silver982K, we employed a reward model trained to optimize spoof-specific consistency and rule-based constraints rather than open-ended language fluency, reducing the likelihood of inherited bias.
6.3 Data Privacy and Consent
FaceCoT also respects individual privacy and consent, particularly when using publicly sourced visual data. While the datasets employed (e.g., WFAS) are released under academic or Creative Commons licenses, we designed the release to be cautious and transparent:
  • No Image Redistribution: The FaceCoT release contains only annotations and metadata; no image data is redistributed or exposed.
  • Transparent Documentation: The final dataset release will include: (1) explicit documentation of all source datasets and their associated licenses; (2) clear usage guidelines requiring downstream users to adhere to original dataset terms and ethical standards.
7 Demonstration of result interpretability
7.1 Demonstration of real face
Since FAS is inherently a binary classification problem, explaining why a face is real is as important as explaining why a face is spoof. To this end, we present a representative real face case (Fig 11(a)) to illustrate how the model perceives authenticity. Rather than relying solely on the absence of spoofing cues, the model proactively identifies positive evidence, including natural and proportionate facial structure, realistic skin texture consistent with illumination, and coherence between facial appearance and the surrounding environment. This unified reasoning pipeline ensures that both real and spoof faces are explained under the same framework, with real cases supported by explicit authenticity cues.
7.2 Demonstration of eleven benchmark dataset
@@
(a) An interpretable result case on a real face
(b) An interpretable result case on CASIA-MFSD
(c) An interpretable result case on CASIA-SURF-3DMask